OpenStudy (the_musicisonpoint):

Write an equation for the line that is parallel to the given line and passes through the given point. y = x + 8; (2, 16) A. y = x + 6 B. y = x – 6 C. y = x D. y = x + 14

OpenStudy (dayakar):

the equation of the line parallel to y=x+8 is in the form of y= x+ K { to find k value plug (x ,y) = (2, 16) in the equation] can u do it

OpenStudy (the_musicisonpoint):

16/2+8?

OpenStudy (dayakar):

16 = 2+k 16-2 =k 14 =k therefore y=x+k y= x+14